OpenStudy (anonymous):

Line AB has the equation 3x - 4y + 5 = 0. Point (p, p+2) lies on line AB. What is the value of p?

OpenStudy (shamim):

Put x=p y=p+2 In ur given equation

OpenStudy (shamim):

3p-4(p+2)+5=0 3p-4p-8+5=0 p=?

OpenStudy (shamim):

Response plz!!!!!

rvc (rvc):

@LOUDxPACK u there?